OpenStudy (anonymous):

What is the radius of a circle with a circumference of twenty-three point five five meters? Use three point one four for pi and show your work.

OpenStudy (jdoe0001):

so the circumference is 23.55m, we know that the equation to get the circumference of a circle is \(\boldsymbol{2\pi r}\) \(\bf c = 2\pi r \\ \quad \\ 23.55 = 2\pi r\\ \quad \\ \textit{solving for "r", we get}\\ \quad \\ \cfrac{23.55}{2\pi} = r\)

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

okay thanks..

OpenStudy (jdoe0001):

yw